An object of charge, ?? = 2.0????, and mass, ?? = 1.4??, is shot with an initial horizontal velocity of ??⃗???? = 2.0??104 ???? ??̂ into a parallel plate capacitor on which the upper plate is

held at +12,000V and the lower plate at 0V. If the separation distance between the plates is ?? = 5????, find the vertical distance away from the top plate that the object will emerge after travelling through the capacitor a length ?? = 50.0????. Neglect fringe effects. Include gravitational effects where ??⃗ = −9.8 ???? ??̂.
Answer #1

Time of travel = L/v0x = 0.50m/(2e4 m/s) = 25e-6 s

net acceleration a = acceleration due to electrostatic force + acceleration due to gravity

= qE/m + g

= -2e-6*(12000/5e-3)/1.4e-3 -9.8

= -3438 m/s^2

vertical distance away from top plate = -0.5 at^2

= 0.5*3438*25e-6^2

= 0.00000107 m

= 1.07*10^-6 m answer